### v1

# DAPPER: Discriminability-Aware Policy-to-Policy Preference-Based Reinforcement Learning for Query-Efficient Robot Skill Acquisition

Abstract
Preference-based Reinforcement Learning (PbRL) enables policy learning through simple queries comparing trajectories from a single policy. While human responses to these queries make it possible to learn policies aligned with human preferences, PbRL suffers from low query efficiency, as policy bias limits trajectory diversity and reduces the number of discriminable queries available for learning preferences. This paper identifies preference discriminability, which quantifies how easily a human can judge which trajectory is closer to their ideal behavior, as a key metric for improving query efficiency. To address this, we move beyond comparisons within a single policy and instead generate queries by comparing trajectories from multiple policies, as training them from scratch promotes diversity without policy bias. We propose Discriminability-Aware Policy-to-Policy Preference-Based Efficient Reinforcement Learning (DAPPER), which integrates preference discriminability with trajectory diversification achieved by multiple policies. DAPPER trains new policies from scratch after each reward update and employs a discriminator that learns to estimate preference discriminability, enabling the prioritized sampling of more discriminable queries. During training, it jointly maximizes the preference reward and preference discriminability score, encouraging the discovery of highly rewarding and easily distinguishable policies. Experiments in simulated and real-world legged robot environments demonstrate that DAPPER outperforms previous methods in query efficiency, particularly under challenging preference discriminability conditions.
Index Terms:
Reinforcement Learning, Human Factors and Human-in-the-Loop, Legged Robots.
I Introduction
Preference-based Reinforcement Learning (PbRL) has gained attention as it enables the learning of control policies through simple query responses [1, 2]. In typical PbRL frameworks, two state-action trajectories are sampled from interactions between a single policy and the environment, and a human selects the more preferred one. Based on these preference labels, a reward model that approximates human preferences is learned, and the control policy is updated through reinforcement learning. By iteratively updating the reward model with new query responses and optimizing the policy accordingly, PbRL gradually learns a policy that reflects human preferences [3, 4].
A key challenge in PbRL is that diverse state-action trajectories are required to efficiently learn human preferences via queries [5, 6]. Ideally, each update of the reward model should lead to the acquisition of new policies that can generate novel state-action trajectories. However, when using a single policy, policy bias [7, 8] becomes a critical problem. Policy bias refers to the limited behavioral diversity that arises when learning from a single policy, as new policies are trained as incremental updates of previous ones. Because each policy is a continuation of the last, their behaviors tend to be strongly correlated, making it difficult to explore new actions and fully exploit updated reward functions. As a result, the collected trajectories tend to be similar, and behavioral diversity is significantly limited.
Figure 1: Overview of the proposed learning framework: Multiple policies with diverse feature representations are sequentially learned. Queries are generated by sampling from these policies based on human preference discriminability, ensuring easier labeling. This approach enhances query efficiency by increasing the number of discriminable queries.
To address this issue, we focus on the ease of labeling and introduce preference discriminability as a key metric. This metric quantifies whether a human can determine how much two trajectories in a query are closer to their ideal behavior. When one trajectory is clearly closer to the ideal, labeling is straightforward. However, if both trajectories are similarly distant from the ideal, it becomes difficult for the human to judge, and a preference label may not be provided. By generating queries with high preference discriminability, we present clearer differences between trajectories, making them easier to label and thereby improving query efficiency, as a greater number of discriminable queries can be used to learn the human preference.
Building on this perspective, we focus on moving beyond comparisons within a single policy and instead generate queries by comparing trajectories from different policies as shown in Fig. 1. Training multiple policies encourages the emergence of more diverse behaviors, since learning different policies from scratch is not constrained by the policy bias. This approach avoids the tendency of gradual policy updates to produce similar behaviors and supports the generation of richer and more varied state-action trajectories. However, simply increasing diversity does not guarantee that the resulting queries will be easy for humans to label. To ensure that diversity leads to meaningful and discriminable comparisons, we further introduce a preference discriminability-aware query generation mechanism that estimates the discriminability of trajectory pairs.
To this end, we propose a novel framework, Discriminability-Aware Policy-to-Policy Preference-Based Efficient Reinforcement Learning (DAPPER), which integrates two key ideas: preference discriminability and trajectory diversification achieved by multiple policies. DAPPER approximates preference discriminability as a learnable function by training a discriminator that estimates whether a given query is likely to be distinguishable by humans, based on the history of queries labeled as indistinguishable. During training, DAPPER probabilistically samples queries based on their predicted preference discriminability, promoting the selection of discriminable queries. To enhance trajectory diversity, DAPPER trains multiple policies. Each time the reward model is updated, a new policy is retrained from scratch. Finally, during policy learning, we not only maximize for the reward of human preference but also incorporate the preference discriminability score as an auxiliary reward. This dual-objective approach encourages the discovery of policies that yield both high-reward and easily distinguishable behaviors.
We demonstrated the effectiveness of DAPPER in a legged robot simulator and real-world robot environment. DAPPER achieves successful policy learning with fewer queries than previous methods. Furthermore, compared to previous methods, only DAPPER consistently learn policies under the conditions that preference discriminability becomes increasingly challenging. We also applied the learned policies to real-world environments.
II Related work
II-A Learning Control Policies
This section describes previous methods for learning control policies and the differences from our approach.
Imitation Learning: Learning from human demonstrations has been explored in previous studies [9, 10]. However, creating demonstrations for complex mechanisms such as quadrupedal robots is challenging. While recent efforts have focused on improving demonstration efficiency by refining demonstration methods, these approaches still require demonstrators to have prior knowledge of the policy features being taught. Moreover, effective data collection depends on a clear understanding of the desired policy features beforehand [11]. For non-experts, predicting and defining the precise behavior of a quadrupedal robot in advance is difficult, making it challenging to acquire the intended policy features.
Reinforcement Learning: Reinforcement Learning (RL) enables robots to autonomously acquire arbitrary policies through trial-and-error interactions with the environment by designing a reward function [12, 13]. However, designing an appropriate reward function that results in the desired behavior requires a carefully designed reward, often incorporating multiple auxiliary terms and considerable manual effort. Moreover, reward function design requires expert knowledge, making it difficult for non-experts to define rewards that effectively capture their intended requirements. As a result, acquiring policies resulting in specific motion patterns through RL remains difficult for non-experts.
Preference-based Reinforcement Learning: Potentially, even non-experts can evaluate and distinguish desirable robot behaviors by responding to video-based comparison queries, enabling the learning of policies that align with human preferences [1, 6]. However, PbRL requires collecting a large number of queries, making policy learning infeasible within a practical timeframe. To the best of our knowledge, this paper is the first to apply PbRL to learning quadrupedal robot locomotion. To achieve this, we aim to improve query efficiency, enabling policy learning within a reasonable timeframe.
II-B Improving Query Efficiency in PbRL
This section reviews previous works for improving query efficiency in PbRL and discusses the differences from our work.
Data Augmentation: In this approach, in addition to human annotations, the preference reward model under training is also used to automatically assign labels. Specifically, when a query sample is fed into the model and the predicted probability (ranging from 0 to 1) that one trajectory is preferred over the other shows a clear margin, the model’s prediction is used as the label instead of querying a human. This enables automatic dataset expansion and reduces the need for human responses, thereby improving query efficiency [2, 6].
Robustifying to Incorrect Human Feedback: Incorrect human feedback can mislead the learning process, reducing the effectiveness of each query. To improve query efficiency under such conditions, robust learning techniques are employed to reduce sensitivity to label noise. These include filtering out inconsistent responses and using robust loss functions that reduce the influence of potentially incorrect labels. By mitigating the negative impact of noisy data, these methods enable more reliable policy updates from fewer queries, thereby enhancing overall query efficiency [3, 4].
Uncertainty Sampling: Uncertainty sampling aims to improve query efficiency by prioritizing queries near the decision boundary of the preference reward model, where the output probability is close to $0.5$. These queries are expected to provide the most informative feedback for refining the reward function. By focusing on ambiguous cases, the model can collect targeted training data that reduces uncertainty and sharpens its decision boundaries [5, 6].
Ours: While previous approaches improve query efficiency by leveraging model confidence [2], addressing annotation noise [5], or actively selecting informative queries [6], they do not explicitly account for the labeling difficulty experienced by human annotators. In contrast, our framework improves query efficiency by modeling and leveraging human preference discriminability, which helps reduce indistinguishable queries and enhances query efficiency. This is particularly important in tasks where distinguishing task features is inherently difficult. If a query is not distinguishable to the human, a preference label cannot be obtained.
III Preliminaries
III-A Reinforcement Learning
Reinforcement Learning (RL) assumes a scenario in which an agent takes actions through observed state [14, 15]. The agent receives feedback about its actions in the form of a reward. The goal of the agent is to learn actions to maximize total reward. The standard formalization of a RL problem builds on the notion of a Markov Decision Process (MDP) and consists of: A continuous state space $S\subseteq\mathbb{R}^{n}$ which represents an observation from the environment; A continuous action space $A\subseteq\mathbb{R}^{m}$ representing the set of actions available to the agent; state transition probability $\mathcal{P}:S\times A\rightarrow P(S)$, where $P(S)$ denotes the probability that action $a$ in state $s$ leads the agent to state $s^{\prime}$; $\gamma\in[0,1)$ is the discount factor. The agent interacts with the environment by taking an action $a\in A$ based on the current state $s\in S$, receiving a reward $r(s,a)$, and transitioning to a new state $s^{\prime}\in S$ according to the probability distribution $\mathcal{P}(\cdot\mid s,a)$.
The agent moves through the state space by repeatedly taking actions, thereby generating a trajectory $\tau=(s_{0},a_{0},s_{1},\dots,s_{t},a_{t},s_{t+1})$, where $s_{0}$ is the state the agent starts with, $a_{t}$ is the action it takes in state $s_{t}$, and $s_{t+1}$ is the next state produced by this action.
The most common task in reinforcement learning is to learn a policy $\pi:S\rightarrow A$ that prescribes the optimal action for the agent in each state. More specifically, the objective is often defined as maximizing the total rewards: $V_{\pi}(s)=\mathbb{E}_{\pi,\mathcal{P}}\Bigl{[}\textstyle\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{t})\,\Bigl{|}\,s_{0}=s\Bigr{]}$. Accordingly, common approaches have focused on learning an optimal policy by maximizing the value function $V$. Among various policy optimization algorithms, Proximal Policy Optimization (PPO) is recognized as an effective means to achieve this objective [12].
Figure 2: Learning process of DAPPER: Our framework consists of three steps. (1) Learning current policy $\pi_{i}$ with reward $r$. (2) Query collection. This step makes queries from current policy $\pi_{i}$ and sampled policy from policy dataset $\Pi$ based on discriminability. Then, trajectories of each policy are generated $\tau_{x}$ and $\tau_{i}$, and queries are labeled $y$. (3) Reward Update. This step trains the preference reward model $R^{H}$ and discriminator $D$ from query dataset $\mathcal{D}$.
III-B Preference-based Reinforcement Learning
In Preference-based Reinforcement Learning (PbRL), the agent learns a reward function based on human feedback and then uses this reward function to learn the optimal policies [1]. The process is as follows.
First, the agent collects state-action trajectories $\tau$. This transition is provided as preferences over pairs of trajectories, expressed as $P(\tau_{i}\succ\tau_{j})$, where $P(\tau_{i}\succ\tau_{j})$ represents the probability that trajectory $\tau_{i}$ is preferred over trajectory $\tau_{j}$. Basically, for each trajectory of a query pair, PbRL utilizes an interval clipped from the trajectory to a specific step length.
Human preference is modeled by logistic regression as follows:
$P(\tau_{i}\succ\tau_{j})=\frac{\exp(r(\tau_{i}))}{\exp(r(\tau_{i}))+\exp(r(\tau_{j}))}$ (1)
where $r(\tau)$ is the cumulative reward function for trajectory $\tau$.
Given the feedback data $\{(\tau_{i},\tau_{j})\}_{k=1}^{N}$, the reward function $r(\tau)$ is optimized by minimizing the binary cross-entropy loss function:
$\mathcal{L}$ $=$ $-\textstyle\sum_{k=1}^{N}\big{[}y_{k}\log P(\tau_{i_{k}}\succ\tau_{j_{k}})+$ (2)
$(1-y_{k})\log P(\tau_{j_{k}}\succ\tau_{i_{k}})\big{]}$
where $y_{k}$ is the query label from the human preference, with $y_{k}=1$ if $\tau_{i_{k}}$ is preferred over $\tau_{j_{k}}$, and $y_{k}=0$ otherwise.
Using the estimated reward function $r(\tau)$, the agent learns a policy $\pi(a|s)$ within the reinforcement learning framework. The agent continuously receives human feedback as it takes new actions, updating the reward function $r(\tau)$ iteratively. This iterative process involves re-estimating the reward function based on new feedback data and refining the agent’s policy.
IV Proposed Framework
We propose Discriminability-Aware Policy-to-Policy Preference-Based Efficient Reinforcement Learning (DAPPER), a framework for query-efficient policy learning through easy-to-label queries. An overview is shown in Fig. 2. Queries compare trajectories from different policies. Each query includes a preference label and a discriminability label assessing whether superiority can be determined. To enhance policy distinction, a discriminator is trained with the preference reward model to guide query sampling and promote diverse policies.
The following sections describe the proposed framework. Section IV-A introduces its main components. Using the reward functions described in Section IV-B, DAPPER learns policies that reflect human preferences through the iterative execution of three steps: (1) Policy Learning, (2) Query Collection, and (3) Reward Update. Starting from Section IV-C, the subsequent sections provide details of each learning step. The overall procedure is summarized in the pseudo-code shown in Algorithm 1.
Set parameters described in Table I
Set policy dataset $\Pi$ and query dataset $\mathcal{D}$
while $i=1,2,...,\text{until converge}$ do
       Set current policy $\pi_{i}$ and rollout buffer $\mathcal{B}_{i}$
       while until converge do
             for $e=1,2,...,E$ do
                   for $t=1,2,...,T$ do
                         Get observation $s_{t}$ from environment
                         Take action $a_{t}$ by $\pi_{i}(s_{t})$
                         Caclulate reward $r_{t}$ by Eq. (4)
                         Push $(s_{t},a_{t},r_{t})$ to $\mathcal{B}_{i}$
            Update policy $\pi_{i}$ by RL update scheme (e.g., [13]) using $\mathcal{B}_{i}$
      Store learned policy to policy dataset $\pi_{i}\rightarrow\Pi$
       for $n=1,2,...,N$ do
             Choose policy $\pi_{x}$ by Eq. (6)
             Generate trajecoties $\tau_{x}$, $\tau_{i}$ and conduct queries
             Store label $y$ to query dataset $(\tau_{x},\tau_{i},y)\rightarrow\mathcal{D}$
      Update preference reward model $R^{H}$ by Eq. (7)
       Update discriminator $D$ by Eq. (8)
Algorithm 1 DAPPER
TABLE I: Learning Parameters in Experiments
Para. Meaning Value
$\alpha$ Coefficient of Query Sampling $10^{-3}$
$\beta$ Coefficient of Policy Exploration $0.6$
$\gamma$ Discount factor of RL $0.99$
$N$ Query number per iteration $10$
$T$ Number of steps per episode $128$
$E$ Number of episodes per one policy update cycle (Parallel environment number) $128$
IV-A Framework Components
DAPPER consists of the following components.
Policy $\pi_{i}(s;\theta_{i})$: A policy used for learning human preferences is parameterized by $\theta_{i}$. Multiple policies with distinct parameters $\theta_{i}$ are trained for use in query comparisons, and the resulting policies are collected in the policy dataset $\Pi\ni(\pi_{1},\pi_{2},\dots,\pi_{i},\dots)$.
Trajectory Generator $\mathcal{G}(\pi)$: A function that samples a trajectory $\tau$ from a policy $\pi$. This function is used to make policy-based queries and also serves as input to the discriminator for comparing policies. We define the generator as: $\mathcal{G}(\pi)=\tau\sim\pi,\quad\text{where }\tau=(s_{0},a_{0},s_{1},\dots,a_{T-1},s_{T})$. Here, $T$ denotes the episode length.
Preference Reward Model $R^{H}(s,a;\phi)$: A reward function model for learning human preferences parameterized by $\phi$, defined identically to PbRL as described in Section III-B. Based on this reward function, human preferences are estimated as shown in Eq. (1).
Discriminator $D(\mathcal{G}(\pi_{i}),\mathcal{G}(\pi_{j});\psi)$: A model that estimates the probability of whether preferences of policy pairs are discriminable by humans parameterized by $\psi$. By learning from previously trained policies, it functions as a reward estimation, ensuring that new policy remain distinguishable. Additionally, it is used to prioritize policy sampling when selecting policies for queries.
IV-B Reward Formulation
Our framework proposes policy-based queries, in contrast to previous works that employ trajectory-based queries from a single policy as shown in Eq. (1). Accordingly, we define human preference as follows:
$P(\pi_{i}\succ\pi_{j})=\frac{\exp[R^{H}(\mathcal{G}(\pi_{i}))]}{\exp[R^{H}(\mathcal{G}(\pi_{i}))]+\exp[R^{H}(\mathcal{G}(\pi_{j}))]}.$ (3)
In DAPPER, in addition to aligning with human preferences, the framework aims to learn policies whose preferences can be discriminated in queries. To this end, the following reward function $r$ is defined from preference reward model $R^{H}$ and discriminability reward $R^{D}$, as shown below:
$r=(1-\beta)R^{H}+\beta R^{D}.$ (4)
The hyperparameter $\beta$ trades off alignment with preference reward learned from human queries and policy diversity.
Let $(s_{t},a_{t})$ be a one sample set of current step $t$ obtained from the interaction between the current policy $\pi_{i}$ and the environment, the discriminability reward $r^{D}$ is calculated from the discriminator $D$ as follows:
$R^{D}(s_{t},a_{t})=\frac{1}{N^{\prime}}\sum\nolimits_{\pi_{x}\in\Pi}D((\bar{\tau_{i}},s_{t},a_{t}),\mathcal{G}(\pi_{x})),$ (5)
where $\bar{\tau_{i}}=(s_{0},a_{0},s_{1},\dots,a_{t-1},s_{t-1})$ denotes the partial trajectory of policy $\pi_{i}$ up to step $t$, and $N^{\prime}$ is the number of previously learned policies $(\pi_{1},\pi_{2},\dots,\pi_{i-1})\in\Pi$ used for comparison.
IV-C Policy Learning
In this process, the policy $\pi_{i}$ for the current iteration $i$ is initialized and trained until the reward $r$ defined in Eq. (4) converges. Once training converges, the learned policy is added to the policy dataset $\Pi$ and used for query generation.
IV-D Query Collection
In this process, queries are generated from the learned policies, and labels are collected through the queries. For query generation, DAPPER utilizes the discriminator $D$ to perform prioritized sampling of policies used in queries, aiming to enhance query discriminability. Specifically, one of the policy pairs is selected from the current policy $\pi_{i}$, which has high discriminability against past policies learned through $R^{D}$. The other policy $\pi_{x}$ is selected from the policy dataset $\Pi\ni(\pi_{1},\pi_{2},\cdots,\pi_{i-1})$. At this stage, the discriminability probability $D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{i}))$ between the current policy $\pi_{i}$ and a past policy $\pi_{x}\in\Pi$ is estimated. Based on this probability, policies for queries are sampled according to the sampling probability $p_{x}$ defined as follows:
$p_{x}=\frac{\exp\big{[}\alpha D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{i}))\big{]}}{\sum_{j=1}^{N}\exp\big{[}\alpha D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{j}))\big{]}}$ (6)
where $\alpha$ is a temperature parameter that controls sampling sharpness: higher values of $\alpha$ favor more discriminable policies, while lower values result in more uniform sampling. Sampling is performed without duplication of any policy pairs already present in the query dataset $\mathcal{D}$.
Then, using the sampled policies, a state-action trajectory $\tau$ is generated from the environment, and the label $y=P(\pi_{x}\succ\pi_{i})$ is collected. In this process, the label $y$ takes values from $\{0,0.5,1\}$, where $y=1$ indicates that $\pi_{x}$ is preferred over $\pi_{i}$, $y=0$ indicates that $\pi_{i}$ is preferred over $\pi_{x}$, and $y=0.5$ signifies that both policies are considered equally preferable or that ranking them is difficult. The obtained labeled data $(\pi_{x},\pi_{i},y)$ is added to the query dataset $\mathcal{D}$.
IV-E Reward Update
IV-E1 Execution Flow
Using the query dataset $\mathcal{D}$, the preference reward model $R^{H}$ for calculating $r^{H}$ and the discriminator $D$ for calculating $r^{D}$ are updated. After the models $R^{H}$ and $D$ have converged, they are utilized in training the next iteration’s policy $\pi_{i+1}$. This process enables learning policies that reflect human preferences from queries while also promoting the acquisition of policies that remain distinguishable from past policies $\Pi$.
IV-E2 Preference Reward
Using the separable label dataset $y\in\{0,1\}$ from the query dataset $\mathcal{D}$, the preference reward model $R^{H}$ is updated. The loss function is defined as follows:
$\mathcal{L}_{\phi}$ $=$ $-\mathbb{E}_{(\pi_{i},\pi_{x},y)\sim\mathcal{D}}\Big{[}y\log P(\pi_{x}\succ\pi_{i})$ (7)
$+(1-y)\log P(\pi_{i}\succ\pi_{x})\Big{]}.$
IV-E3 Discriminability Reward
Using the dataset with inseparable labels $y\in\{0.5,\text{other}\}$ from the query dataset $\mathcal{D}$, the discriminator $D$ is updated. The corresponding loss function is defined as follows:
$\mathcal{L}_{\psi}$ $=$ $-\mathbb{E}_{(\pi_{i},\pi_{x},y)\sim\mathcal{D}}\Big{[}(1-\delta(y))\log D(\mathcal{G}(\pi_{x}),\mathcal{G}(\pi_{i}))$ (8)
$+\delta(y)\log D(\mathcal{G}(\pi_{i}),\mathcal{G}(\pi_{x}))\Big{]}$
where $\delta$ is a delta function indicating discriminability, defined as $\delta=1\text{ if }y=0.5,\;\delta=0\text{ if }y=0\ \text{or}\ 1$.
TABLE II: Constraint parameters: Following the definitions in Section V-A2, the constraint target parameters for each task are defined. These parameters are estimated by feature extractor $f$ from state-action trajectories $\tau$. The threshold is set to a common value of $d=10$ for all parameters.
Parameter Target: $x^{*}$ Marge: $g$
Body Velocity of X Direction [m/s] 1 0.2
Body Velocity of Y Direction [m/s] 0 0.2
Body Rotation of Z Axis [degree] 0 10
Collisions on Bodies and Legs 0 0
V Experiments
The following experiments investigate whether the proposed framework can achieve higher query efficiency than existing methods by considering preference discriminability. In comprehensive comparative experiments, it is impractical for an actual human to manually label all queries due to the enormous time and human cost. To address this, we define a simulated annotator model and perform automatic labeling. We first formulate the simulated annotator’s preference model and preference discriminability as mathematical expressions. Subsequently, we conduct a human subject experiment to validate the simulated annotator model by examining whether actual human responses to trajectory comparison queries can be explained by the preference discriminability defined in the model (at Section V-C). From the experimental results, we quantitatively determine an approximate threshold for preference discriminability and define multiple levels around it. These thresholds are then used in the subsequent PbRL experiments (from Section V-D) to automate query labeling with the simulated annotator model, enabling efficient evaluation of the proposed framework.
In our experiments, we aim to address the following research questions:
  • To what extent does the preference discriminability defined in the simulated annotator model explain actual human responses to trajectory comparison queries? (Section V-C)
  • Does our framework improve query efficiency compared to previous methods? (Section V-D)
  • How does hyperparameter selection affect query efficiency and preference discriminability in our framework? (Section V-E)
  • Can the proposed method maintain its effectiveness as the number of feature dimensions that parametrize the target robot behaviors increases? (Section V-F)
  • Is our framework applicable to sim-to-real policy transfer? (Section V-G)
TABLE III: Terms of feature extraction: The second and third columns show the parameter ranges, while the fourth column presents the target feature values $f^{*}$. The remaining columns indicate the walking patterns and the features used in each pattern, where a checkmark means the feature follows $f^{*}$, and a numerical value means it follows the specified value. “Gait Cycle” and “Phase Difference” are calculated from the timing of foot-ground contact. “FL” (Front Left), “FR” (Front Right), “RL” (Rear Left), and “RR” (Rear Right) refer to the respective legs of a quadrupedal robot. “Gait Cycle” is the duration of a foot’s contact cycle, while “Phase Difference” denotes the delay between two legs’ contact timings. Both are updated at each foot contact, with the most recent value used at step $t$. “Phase Difference” is given as the delay ratio to the current gait cycle. “Body Height” and “Body Incline Angle” are computed at each step $t$.
Parameter Min Max Target $f^{*}$ Posture Trot Only Normal Crawl
Body Height [cm] 25 65 45 45 - 55 30
Body Incline Angle [degree] -10 10 0 -
Gait Cycle (FL) [s] 0.2 0.4 0.5 -
Phase Difference (FL-FR) [%] 0 100 50 -
Phase Difference (FL-RL) [%] 0 100 50 -
Phase Difference (FL-RR) [%] 0 100 0 -
V-A Common Settings
[…]
V-A1 Feature Extraction
To quantify the degree of alignment, we utilize the policy feature representation $f(s,a)$ extracted by a feature extractor introduced in previous PbRL works [19, 20]. In the feature space defined by $f$, we specify a target feature $f^{*}$ that represents the desired behavior in accordance with human preferences. The target features are listed in Table III. In this paper, feature extraction is also performed when comparing different policies. In such cases, trajectory $\tau$ is generated by a policy $\pi$, then the average feature representation is calculated as $f(\pi)=\frac{1}{T}\textstyle\sum_{(s,a)\in\tau}f(s,a)$, where $T$ denotes the episode length.
V-A2 Constraint Terms of LPPO
This paper adopts LPPO to learn a policy that satisfies constraints while suppressing unrealistic behaviors that may hinder accurate query labeling. Constraint term is defined by indicator function $\mathbb{I}$, allowing the cost function $c$ to be expressed as follows: $c(s_{t},a_{t})=\mathbb{I}\left[|x_{t}-x^{*}|>g\right]$, where $x_{t}$ denotes a current value of constraint target parameter, $x^{*}$ is the target value, and $g$ is the permissible margin of error. This formulation penalizes each timestep in which the deviation from the target exceeds the margin $g$. To enforce the constraint over an episode of length $T$, LPPO defines an threshold $d$ on the total number of penalty-inducing steps: $\sum_{t=1}^{T}c(s_{t},a_{t})\leq d$. This constraint ensures that the number of steps in which the system deviates significantly from the desired behavior remains within an acceptable bound. The constraint terms are listed in Table II.
V-A3 Labeling Queries from Simulated Annotator Model
To reduce the human cost associated with large-scale evaluation, we employ a simulated annotator model. Specifically, we define a simulated annotator’s preference model and preference discriminability as mathematical expressions, enabling automatic labeling. The simulated annotator assigns labels $y\in{0,0.5,1}$ based on the extracted policy features $f(\pi)$ and a predefined target feature $f^{*}$. All feature values are normalized to the range $[0,1]$ according to the parameter specifications listed in Table III.
To estimate preference and discriminability labels, we define a unified discriminability metric based on the difference in feature errors between two trajectories $\tau_{i}$ and $\tau_{j}$ with respect to a target feature $f^{*}$. Specifically, we use the following scale-normalized L1-based distance:
$d^{\text{pref}}(\tau,f^{*})=\frac{1}{\sqrt{|f|}}\textstyle\sum_{k=1}^{|f|}|f_{k}(\tau)-f^{*}_{k}|,$ (9a)
$d^{\text{disc}}(\tau_{i},\tau_{j},f^{*})=|d^{\text{pref}}(f(\tau_{i}),f^{*})-d^{\text{pref}}(f(\tau_{j}),f^{*})|,$ (9b)
where $|f|$ denotes the number of feature dimensions. Here, the trajectory with the smaller feature error $d^{\text{pref}}$ is labeled as more preferred ($y\in\{0,1\}$), while the discriminability label ($y\in\{0.5,\text{other}\}$) is determined by whether preference discriminability $d^{\text{disc}}$ exceeds a predefined threshold. If the value is above the threshold, the query is considered distinguishable; otherwise, it is treated as indistinguishable.
[…]
V-B Comparing Methods
We compared the following methods to verify the effectiveness of the proposed method:
  • Baseline [1]: A method that generates queries from pairs of state-action trajectories obtained from different episodes using a single policy.
  • SURF [2]: Same as the Baseline, queries are generated from a single policy. However, this method automatically labels queries when the preference values calculated from the learned reward function are sufficiently high, thereby increasing the number of training samples.
We also verified the following two ablation methods of our framework:
  • Ours w/o $R^{D}$: Same as the proposed framework, queries are generated from pairs of state-action trajectories obtained from different policies. However, the reward function $R^{D}$, which encourages acquiring policies with distinct features, is not used.
  • Ours with Distance Maximization (Ours-DM): This is the proposed framework with different $R^{D}$ which is calculated based on the feature distance from previously learned policies using the following equation: ${R^{D}}^{\prime}=\frac{-1}{N^{\prime}}\sum_{\pi_{x}\sim\Pi}|f(\pi_{i})-f(\pi_{x})|$. This setting is included to evaluate whether simply maximizing the distance between policies is more effective than Ours.
V-C Analysis of Actual-Human Preference Discriminability
V-C1 Settings
This paper proposes a framework for generating queries that are easier for humans to distinguish, based on preference discriminability. This section examines whether actual humans can distinguish similar queries and how their ability declines with increasing similarity. To control for individual preference variability, we define a target feature $f^{*}$ and treat the policy closer to $f^{*}$ as the correct answer. We then analyze the proportion of indistinguishable responses ($y=0.5$) as a function of $d^{\text{disc}}$ to quantify actual-human preferecne discriminability. Additionally, we evaluate the accuracy of query labels when participants express a clear preference ($y\neq 0.5$). This experiment uses the “Posture” feature values defined in Table III.
V-C2 Results
[…]
Figure 4: Evaluation of query efficiency: (a) Comparison to previous works. Top: Discrimination rate of queries. Middle: Converge query number for reaching the feature error threshold ($d^{\text{pref}}<0.02$), empirically defined by the convergence values of each method. If not converged within $2\times 10^{3}$ queries, the value is clipped at $2\times 10^{3}$. The threshold is shown as a red dashed line in the bottom plot. Bottom: Minimum values of feature error $d^{\text{pref}}$ achieved by the learned policies. “Small”, “Medium”, and “Large” denote levels of preference discriminability $d^{\text{disc}}$. Each curve shows the mean and variance over five experiments. (b) Learned policy features per query number. This experiments use $d^{\text{disc}}$ = “Medium”. The target feature $f^{*}$ is located at the intersection of the red dashed lines. The number of plotted points corresponds to the queries collected until the feature error falls below the threshold ($d^{\text{pref}}<0.02$), indicated by the red frame.
Figure 5: Performance comparison of DAPPER components: (Left) different temperature patterns of query sampling. (Right) different scaling coefficient patterns of policy exploration. “Small”, “Medium”, and “Large” denote levels of preference discriminability $d^{\text{disc}}$. Each curve plots the mean and variance over five experiments.
V-D Comparison of Query Efficiency with Other Methods
V-D1 Settings
The proposed framework improves query efficiency by utilizing queries from multiple policy trajectories instead of a single policy trajectory, thereby enhancing query discriminability to improve query efficiency. To verify its effectiveness, we compare the query discrimination rate and the number of queries required to reach the predefined target feature $f^{*}$ (as shown in Table III) against the other methods described in Section V-B. We evaluate all methods across three levels of preference discriminability $d^{\text{disc}}$ to examine their robustness under increasingly difficult query discrimination conditions. This experiment uses target feature values of “Posture” as defined in Table III.
V-D2 Results
As shown in Fig. 4(a), Ours (DAPPER) achieves the target features with the fewest number of queries. When queries become harder to distinguish due to increased preference discriminability $d^{\text{disc}}$, the number of required queries increases. In particular, when $d^{\text{disc}}$ is “Large”, Ours requires the fewest queries; by contrast, the ablation variant requires over three times more queries, and the trajectory-based method fails to learn the target features within the given number of queries. Ours consistently achieves the highest discrimination rate, and under the “Large” condition, only Ours and Ours-DM are able to generate a sufficient number of discriminable queries for learning a policy that closes the target features.
As shown in Fig. 4(b), Baseline suffers from policy bias, where exploration is restricted by features of previously learned policies, resulting in slow progress toward the target features. In contrast, Ours retrains policies from scratch at each reward model update, avoiding policy bias and rapidly acquiring new features. Consequently, Ours outperforms the Baseline in terms of query efficiency.
V-E Influence of Different Query Exploration/Sampling
V-E1 Settings
As shown in Eq. (4), the reward in DAPPER is adjusted by setting the ratio $\beta$ between $r^{D}$ and $r^{H}$ to balance policy discriminability and convergence toward the target feature values. Additionally, the policy sampling from the policy dataset $\Pi$ for query generation also affects discriminability performance (Eq. (6)). Therefore, this section comprehensively compares these parameters and evaluates their influence on convergence to target features. This experiment uses target feature values of “Posture” as defined in Table III.
V-E2 Results
The results are shown in Fig. 5. Regarding query sampling, a higher temperature coefficient $\alpha$ increases the probability of selecting policies with high discriminability, thereby improving the overall discriminability rate. However, a large $\alpha$ results in biased sampling, as only easily distinguishable policies are selected by the discriminator $D$. Moderately reducing $\alpha$ improves query efficiency by encouraging diverse queries, which better support learning the preference reward $R^{H}$. Regarding policy exploration, a higher weight $\beta$ promotes broader exploration, which is expected to facilitate reaching the optimal policy. However, excessive exploration degrades convergence, as also observed in the results of Ours-DM in Section V-D. Since $\alpha=10^{-3}$ and $\beta=0.6$ consistently yield good performance across different $d^{\text{disc}}$ settings, we use them in all experiments.
Figure 6: Comparison of query efficiency in various tasks: “Trot only” means using feature extraction of trot patterns, “Normal” adds feature extraction of body posture as shown in Table III. This experiments use preference discriminability $d^{\text{disc}}$ = “Medium”, meaning that if a policy’s features are closer than a specified threshold, the policy is considered non-identifiable ($y=0.5$). Each curve plots the mean and variance over five experiments.
Figure 7: Real-world execution of policies trained in simulation: The “Crawl” walking pattern listed in Table III is trained in simulation, and the learned policies were deployed on the ANYmal quadrupedal robot.
V-F Scalability with Respect to Feature Dimensions
V-F1 Settings
This section evaluates the applicability of the proposed framework to different tasks by testing its effectiveness on two additional task settings with a greater number of feature parameters. The experiment uses the target feature values of “Trot Only” and “Normal”, as defined in Table III. In these settings, the number of features required for learning from queries is increased to four and six, respectively, compared to two in the tasks used in the previous section, making the learning process more challenging. All experiments were conducted under the “Medium” preference discriminability condition.
V-F2 Results
As shown in Fig. 6, Ours achieves faster convergence to the target feature $f^{*}$ across all tasks. As the number of features $|f|$ increases, the difficulty of discrimination also rises, and the number of required queries tends to increase accordingly. In particular, previous methods struggle significantly with the “Normal” setting, which requires learning six features, resulting in very few identifiable queries. While the learning difficulty may ultimately depend on how close the policy needs to get to the target feature $f^{*}$, it is possible that even our proposed framework may face limitations as the feature dimensionality increases further. Investigating this potential limitation remains an important direction for future work.
V-G Real-World Application
In this section, we apply the learned control policies from simulation to real-world environments using a sim-to-real transfer approach [16]. The transfer policy is trained in a domain-randomized environment using the preference reward model $R^{H}$ obtained from human queries. The randomized parameters follow the same settings as in previous work [16]. As shown in Fig. 7, the learned policies were applied in the real world. For details of the robot motion, please refer to the supplementary videos.
VI Discussion
This paper focuses on preference discriminability and proposes a robust learning framework that performs effectively even under high discriminability thresholds (i.e., “Large”). However, in practice, some queries that fall within these thresholds may not be labeled as indistinguishable by human annotators but instead receive incorrect superiority labels as shown in Fig. 3(c). This type of labeling noise has been addressed in several prior works [3, 4]. Integrating such noise-robust methods with our proposed approach presents a promising direction for future research.
Some consideration is necessary when defining the terms of feature extraction [20]. In particular, if multiple similar features are selected and their optimization objectives happen to conflict, it may hinder policy learning. For example, if satisfying different features requires the robot to exhibit opposing behaviors, achieving the target may become difficult or even infeasible. While such cases are unlikely in well-structured tasks, they remain a possibility, and attention to feature design can help avoid such issues.
In our framework, policy constraints (Table II) are used to enforce basic requirements such as movement speed or safety, rather than to represent human preferences. These constraints do not need to be strictly maximized, but only sufficiently satisfied within a margin. This approach is particularly effective because such low-level objectives are easy to define and help improve query discriminability. In contrast, high-level preferences such as motion style are difficult to specify explicitly, and are therefore learned through PbRL. This separation allows the system to delegate well-defined tasks to constraints, while using PbRL to capture complex, preference-based behaviors.
VII Conclusion
This paper proposed DAPPER, a framework that improves query efficiency by incorporating human preference discriminability and trajectory diversification achieved by multiple policies. DAPPER trains multiple policies and learns a discriminator to generate queries that are easier for humans to label, addressing the limited diversity problem caused by policy bias inherent in previous trajectory-based approaches. Experiments on simulated and real-world legged robot tasks demonstrated that DAPPER significantly improves query efficiency compared to previous methods, particularly under challenging discriminability conditions.